### v1

@@ Abstract
Deepfakes have recently raised significant trust issues and security concerns among the public. Compared to CNN face forgery detectors, ViT-based methods take advantage of the expressivity of transformers, achieving superior detection performance. However, these approaches still exhibit the following limitations: (1). Fully fine-tuning ViT-based models from ImageNet weights demands substantial computational and storage resources; (2). ViT-based methods struggle to capture local forgery clues, leading to model bias and limited generalizability. To tackle these challenges, this work introduces Mixture-of-Experts modules for Face Forgery Detection (MoE-FFD), a generalized yet parameter-efficient ViT-based approach. MoE-FFD only updates lightweight Low-Rank Adaptation (LoRA) and Adapter layers while keeping the ViT backbone frozen, thereby achieving parameter-efficient training. Moreover, MoE-FFD leverages the expressivity of transformers and local priors of CNNs to simultaneously extract global and local forgery clues. Additionally, novel MoE modules are designed to scale the model’s capacity and select optimal forgery experts, further enhancing forgery detection performance. The proposed MoE learning scheme can be seamlessly adapted to various transformer backbones in a plug-and-play manner. Extensive experimental results demonstrate that the proposed method achieves state-of-the-art face forgery detection performance with reduced parameter overhead. The code will be released upon acceptance.
Keywords:
Deepfakes face forgery detection Mixture-of-Experts generalizability robustness parameter-efficient training.
@@ 1 Introduction
With rapid advancements in AIGC, forged face contents are becoming increasingly sophisticated. The human naked eye can hardly discriminate between fake faces and real ones. Even non-experts can readily implement face manipulation algorithms to produce falsified facial images/videos (i.e., Deepfakes) with a high level of realism. Consequently, the rapid proliferation of Deepfake content on social media platforms has caused pressing security issues, such as disinformation, fraud, and impersonation [29]. Thus, developing accurate and robust face forgery detectors to counter malicious attacks poses a grand challenge.
@@
With the advent of the vision transformer (ViT) [13], ViT architectures have achieved significant success in various computer vision tasks due to their long-range interactions and outstanding expressivity. In the realm of Deepfake detection, numerous ViT-based approaches [12, 55, 73, 67, 28, 15, 46] have been proposed, achieving enhanced accuracy and generalizability. Nevertheless, ViT-based forgery detection approaches still face certain limitations. Firstly, fully training ViT-based models from the ImageNet weights demands substantial computational resources, hindering their deployment or fine-tuning in real-world applications, especially on mobile devices with limited processing power. Secondly, while ViT-based methods exhibit outstanding expressivity, they may struggle to capture forgery features in local abnormal regions, resulting in model bias and limited generalizability.
@@
Compared with previous methods that directly fine-tune the entire ViT parameters, the proposed fine-tuning strategy effectively preserves abundant knowledge from ImageNet and enables the model to adaptively learn forgery-specific features. Additionally, the designed LoRA layers model the long-range interactions within input faces, while the Convpass Adapter layers effectively highlight local forgery anomalies. To this end, the integration of the designed LoRA and Adapter layers leverages the expressivity of transformers and the local forgery priors of CNNs, leading to enhanced generalizability and robustness. Moreover, we design novel MoE modules in both LoRA and Adapter layers to scale the model capacity using fixed activated parameters. MoE aims to dynamically select optimal LoRA and Adapter experts for input faces, further empowering the model’s forgery detection performance.
@@
  • We innovatively integrate LoRA and Adapter with the ViT backbone for face forgery detection. This design simultaneously mines global and local forgery clues in a more efficient manner.
  • We design novel Mixture-of-Experts (MoE) modules to scale the model capacity, which dynamically select optimal forgery experts for input faces. Moreover, the designed MoE modules can be seamlessly adapted to other transformer architectures and boost the detection performance in a plug-and-play fashion.
  • We conduct experiments on six Deepfake datasets and various perturbations. Extensive experimental results demonstrate that MoE-FFD achieves state-of-the-art generalizability and robustness.
@@ 2.1 Face Forgery Detection
Early attempts at face forgery detection primarily relied on extracting handcrafted features, such as the lack of eye blinking [37], inconsistency of head pose [70], face warping artifacts [38], and heart rate anomalies [8]. However, these methods suffer from limited accuracy due to their narrow focus on specific statistical information. In response, learning-based approaches have emerged, leveraging generic network architectures like Xception Network [7], EfficientNet [63], and Capsule Network [49] for forgery feature extraction. Nonetheless, CNN-based methods are prone to overfitting to the training data, resulting in limited generalizability and robustness. Follow-up works, such as Face X-ray [36], F3Net [51], SBI [59], DCL [61], and RECCE [6], have introduced comprehensive forgery frameworks and robust feature extraction techniques, enhancing the model’s generalization capability. With the explosive development of ViT [13], numerous ViT-based methods [12, 55, 73, 67, 28, 15, 46] have been proposed to tackle the Deepfake problem. These approaches take advantage of non-inductive bias and global context understanding, achieving superior face forgery detection performance. However, most of these methods fine-tune the entire ViT model on Deepfake datasets, which is computationally expensive and may lead to loss of valuable knowledge from the ImageNet dataset [9]. In contrast, our approach finetunes external lightweight LoRA and Adapter parameters during training, keeping the ViT backbone fixed with ImageNet weights, thus enabling a parameter-efficient face forgery detection.
@@ 2.2 Parameter Efficient Fine-Tuning (PEFT)
Over the past few years, the size of deep learning models has exponentially increased, especially after the advent of transformers. Consequently, numerous Parameter Efficient Fine-Tuning (PEFT) methods have been proposed to reduce the computational and storage overhead. PEFT only updates a small portion of the model parameters while freezing the majority of pretrained weights. Adapter [18, 62] is a typical PEFT method, comprising a down-sample layer and an up-sample layer, generally integrated into transformer layers and blocks. Low-Rank Adaptation (LoRA) [19] layer aims at updating two low-rank matrices, significantly reducing the trainable parameters. Additionally, Visual Prompt tuning (VPT) [22] augments inputs with extra learnable tokens, which can be regarded as learnable pixels to vision transformers. Scale and Shift Feature Modulation (SSF) [40] advances further by introducing to scale and shift parameters to modulate visual features during training, achieving comparable performance compared with full fine-tuning. Convpass [24] integrates convolutional bypasses into large ViT models, leveraging local priors to improve image classification performance. However, the application of PEFT methods to face forgery detection remains largely unexplored. In this work, we incorporate dedicated forgery Adapter and LoRA layers to mine global and local clues and enhance the detection performance.
@@ 2.3 Mixture of Experts (MoE)
Mixture-of-Experts (MoE) [21] models aim to augment the model’s capacity without increasing computational expenses. MoE comprises multiple sub-experts and incorporates a gating mechanism to dynamically select the most relevant Top-k experts, thereby optimizing the results. The concept of MoE has been widely used in both Computer Vision [52, 41, 48, 57] and Natural Language Processing [56, 31, 4]. Sparse MoE [56] introduces a router to select a subset of experts, ensuring that the inference time is on par with the standalone counterpart. Subsequent methods such as [17, 32, 53] seek to design novel gating mechanisms to enhance the performance on specific tasks. Follow-up works [17, 30, 44] propose leveraging multi-task learning to guide the model to select optimal experts for a given input query. Moreover, some studies apply MoE architectures to domain adaptation [16] and domain generalization [33] tasks. In our work, we draw inspiration from the concept of MoE to automate the selection of different LoRA and Adapter modules for various testing data. The designed MoE modules smartly select optimal experts and outperform previous detection methods by a clear margin.
3 Methods
@@ 3.1.1 Vision Transformer
consists of multiple blocks of Self-Attention and Multi-Layer Perceptron (MLP). For a given input sequence $x\in\mathbb{R}^{N_{t}\times D}$, $x$ is firstly projected to queries $Q\in\mathbb{R}^{N_{t}\times dim}$, keys $K\in\mathbb{R}^{N_{t}\times dim}$, and values $V\in\mathbb{R}^{N_{t}\times dim}$ using three learnable matrices $W_{q}\in\mathbb{R}^{D\times dim}$, $W_{k}\in\mathbb{R}^{D\times dim}$, and $W_{v}\in\mathbb{R}^{D\times dim}$, where $dim$ denotes the hidden dimension:
@@
3.1.2 LoRA
is one popular parameter-efficient tuning method. For a pretrained ImageNet weight matrix $W\in\mathbb{R}^{D\times dim}$, LoRA freezes $W$ during training while adding a product of two trainable low-rank matrices $W^{down}W^{up}$, where $W^{down}\in\mathbb{R}^{D\times r}$, $W^{up}\in\mathbb{R}^{r\times dim}$, and rank $r<<min(D,dim)$. As such, the forward pass is modified as:
$h=x(W+\Delta W)=xW+xW^{down}W^{up}.$ (3)
Compared with ViT, the proposed LoRA significantly reduces the trainable parameter number.
@@ 3.1.3 Convpass Adapter
integrates local priors for visual tasks, constructing Convolutional Bypasses (Convpass) within the Vision Transformer (ViT) framework as adaptation modules. The Convpass Adapter can be formulated as:
@@
where the parameters in MLP are fixed with ImageNet weights during training. Convpass generally consists of several trainable convolutional layers.
Figure 2: Overview of the designed MoE-FFD framework. (a). Overall model structure; (b). Details of MoE-FFD transformer block; (c). Details of the designed MoE Adapter layer; (d). Details of each Adapter expert; (e). Details of the designed MoE LoRA layer; (f). Details of each LoRA expert.
@@ 3.2 Overview of the Proposed Framework
Fig. 2 illustrates the designed MoE-FFD framework. Fig. 2 (a) depicts the ViT backbone, which is initialized with the ImageNet weights. Fig. 2 (b) presents more details regarding the designed transformer blocks. We integrate the designed MoE Adapter layer and MoE LoRA layer with the vanilla ViT blocks.
@@
Similarly, the attention weights $W_{q,k,v}$ in the MoE LoRA are fixed to the ImageNet weights. As shown in Fig. 2 (e), the MoE LoRA module consists of one gate $G_{L}(\cdot)$ and $N$ experts, each with a unique rank. Fig. 2 (f) depicts the LoRA structure, where $\otimes$ indicates matrix multiplication. The output feature is derived by multiplying the input with two learnable low-rank matrices.
@@
where $L_{ce}$ represents the cross-entropy loss. Following [56, 5], we further apply an additional loss $L_{moe}$ to encourage all experts to have equal importance. Further details on $L_{moe}$ will be dedicated at the end of this section.
@@ 3.3 Details of the MoE
Figure 3: Illustration of the designed Convpass Adapter experts.
3.3.2 MoE Adapter Layer.
$\small x_{out}={\rm MLP}(x_{in})+\sum_{j}^{M}G_{A}(x_{in})_{j}E_{A}(x_{in})_{j},$ (8)
$\rm Conv_{1\times 1}^{down}$ and $\rm Conv_{1\times 1}^{up}$ are two $1\times 1$ convolution layers, which down-sample and up-sample the channels, respectively. $\rm Conv_{3\times 3}^{j}$ indicates the specific convolution layer in different experts. We develop $M=5$ types of convolution. These convolutions aim to model different local interactions, facilitating the model to capture abundant local forgery features in input faces.
Fig. 3 shows the designed five convolutions, including (a). Vanilla Convolution, (b). Angular Difference Convolution (ADC), (c). Central Difference Convolution (CDC), (d). Radial Difference Convolution (RDC), and (e). Second-Order Convolution (SOC). The red arrow $\rightarrow$ indicates the subtraction operation. We first formulate the vanilla convolution as (we omit the bias for concision):
$y=\sum_{p\in\Omega_{1}}w_{p}x_{p},$ (10)
3.3.3 Gating Network.
$G(x)={\rm Softmax}({\rm Topk}(H(x),k)).$ (13)
$L_{moe}=CV(Importance(X))^{2};~{}Importance(X)=\sum_{x\in X}G(x),$ (14)
where $CV(\cdot)$ indicates the Coefficient of Variation.
4 Experiments
4.1 Implementation Details
We apply the popular MTCNN face detector [71] to crop the face regions. The proposed framework is implemented on the Pytorch [50] platform. The model is trained using Adam optimizer [25] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate of gating network and other trainable parameters as 1e-4 and 3e-5, respectively. The loss weigh $\lambda$ in Eq. (5) is set as 1. We train the model for 20 epochs on one single 3090 GPU with batch size 32. We follow the data split strategy in [42] for fair comparison.
4.2 Datasets and Evaluation Metrics
The experiments encompass the following six datasets for training and testing: FaceForensics++ (FF++) [54], CelebDF-v2 (CDF) [39], WildDeepfake (WDF) [74], DeepFake Detection Challenge (DFDC) [10], DeepFakeDetection (DFD) [1], and DeepForensics-1.0 (DFR) [23]. FF++ is a widely used dataset in face forgery detection, which includes four face manipulation types: Deepfakes (DF) [2], Face2Face (FF) [65], FaceSwap (FS) [3], and NeuralTextures (NT) [64]. The remaining datasets represent recent five Deepfake datasets with diverse environment variables and various video qualities. We perform cross-dataset and cross-manipulation evaluations to examine the models’ generalization capability. Additionally, we measure the model’s robustness by applying it to perturbed face images with various perturbations and severity levels. Consistent with prior arts, we adopt Area under the ROC Curve (AUC) and Equal Error Rate (EER) as evaluation metrics.
Table 1: Cross-dataset evaluation on five unseen datasets. ‘*’ indicates the trained model provided by the authors. ’†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. #Params indicates the activated parameter number during training.
Method Venue #Params CDF WDF DFDC DFD DFR Avgerage
@@ 4.2 Datasets and Evaluation Metrics
4.3 Comparison with State-of-the-Arts
Cross-Dataset Evaluations. In practical scenarios, the trained models always suffer significant performance drops when tested on unseen datasets. Herein, we conduct cross-dataset experiments to evaluate the models’ generalization capability. We train all models on the FF++ (C23) dataset and directly evaluate them on five unseen Deepfake datasets: CDF, WDF, DFDC, DFD, and DFR. Table 1 presents the frame-level and video-level detection results, divided into top and bottom sections. Video-level detection results are obtained by averaging all frame scores within each video. We bold the best results and underline the second-best results. Notably, MoE-FFD achieves the highest detection performance on four datasets in both frame-level and video-level detection. Additionally, MoE-FFD’s performance is on par with CFM on the DFD dataset. Moreover, MoE-FFD outperforms CFM by a clear margin in average performance. The average frame-level AUC improves by 2.86%, rising from 83.94% to 86.78%. Similarly, MoE-FFD also outperforms others in video-level detection on average.
Cross-Manipulation Evaluations. Existing face forgery detectors are vulnerable to unforeseen manipulation types. To evaluate the generalizability to new manipulation techniques, we perform cross-manipulation experiments on four forgery techniques: Deepfakes (DF), Face2Face (FF), FaceSwap (FS), and NeuralTextures (NT). Table 4.3 presents the results, where models trained on three manipulation types are tested on the remaining one. MoE-FFD exhibits the best average detection results. Compared to the ViT-B baseline, the proposed method achieves an average AUC enhancement of 18.5%, going from 62.3% to 80.8%. In Table 4.3, we further examine models trained on one manipulation type and tested across the other three. AVG denotes the average AUC score across three cross-manipulation trials. MoE-FFD outperforms the baseline in both intra- and cross-manipulation evaluations. Remarkably, MoE-FFD achieves around 99% AUC scores in each intra-manipulation evaluation, highlighting its outstanding detection accuracy in intra-domain evaluations. When compared to the state-of-the-art RECCE method, our MoE-FFD consistently achieves superior average cross-manipulation results across all four evaluation settings. These findings suggest that MoE-FFD is generalized to unseen manipulations.
Table 2: Cross-manipulation detection AUC on the unseen manipulation method (C23).

### v2

@@ Abstract
Deepfakes have recently raised significant trust issues and security concerns among the public. Compared to CNN face forgery detectors, ViT-based methods take advantage of the expressivity of transformers, achieving superior detection performance. However, these approaches still exhibit the following limitations: (1) Fully fine-tuning ViT-based models from ImageNet weights demands substantial computational and storage resources; (2) ViT-based methods struggle to capture local forgery clues, leading to model bias; (3) These methods limit their scope on only one or few face forgery features, resulting in limited generalizability. To tackle these challenges, this work introduces Mixture-of-Experts modules for Face Forgery Detection (MoE-FFD), a generalized yet parameter-efficient ViT-based approach. MoE-FFD only updates lightweight Low-Rank Adaptation (LoRA) and Adapter layers while keeping the ViT backbone frozen, thereby achieving parameter-efficient training. Moreover, MoE-FFD leverages the expressivity of transformers and local priors of CNNs to simultaneously extract global and local forgery clues. Additionally, novel MoE modules are designed to scale the model’s capacity and smartly select optimal forgery experts, further enhancing forgery detection performance. Our proposed learning scheme can be seamlessly adapted to various transformer backbones in a plug-and-play manner. Extensive experimental results demonstrate that the proposed method achieves state-of-the-art face forgery detection performance with significantly reduced parameter overhead. The code is released at: https://github.com/LoveSiameseCat/MoE-FFD.
Index Terms:
Deepfakes, Face Forgery Detection, Mixture-of-Experts, Generalizability, Robustness, Parameter-Efficient Training.
@@ 1 Introduction
With rapid advancements in Artificial Intelligence-Generated Content (AIGC), forged facial content has become increasingly sophisticated, making it difficult for the human eye to distinguish between fake and real faces. Non-experts can easily use face manipulation algorithms to create highly realistic falsified facial images and videos, known as Deepfakes. Consequently, the rapid proliferation of Deepfake content on social media platforms has led to significant security issues, including disinformation, fraud, and impersonation [1]. Even worse, the complexity of deployment environments in real-world applications further deteriorates the performance of detection models. Therefore, developing generalized and robust face forgery detectors to counter malicious attacks remains a substantial challenge.
@@
With the advent of the vision transformer (ViT) [19], ViT architectures have achieved significant success in a wide variety of computer vision tasks due to their long-range interactions and outstanding expressivity. In the realm of Deepfake detection, numerous ViT-based approaches [20, 21, 22, 23, 24, 25, 26] have been proposed, achieving enhanced accuracy and generalizability. For instance, Dong et al. [20] designed a ViT to capture the identity inconsistency between inner and outer face region. Zhuang et al. [22] proposed a ViT to mine unsupervised inconsistency-aware features within each frame. Shao et al. [21] introduced a concise yet effective Seq-DeepFake Transformer to predict a sequential vector of facial manipulation operations.
Nevertheless, ViT-based forgery detection approaches still face several limitations. First, fully training ViT-based models from the ImageNet weights demands substantial computational resources, which hinders their deployment or fine-tuning in real-world applications, particularly on mobile devices with limited processing power. Second, although ViT-based methods exhibit outstanding expressivity, they may struggle to capture forgery features in local abnormal regions, resulting in model bias and limited generalizability. Finally, previous methods focus on certain forgery artifacts, but it is challenging to empirically select the optimal features in unpredictable application scenarios.
@@
Compared with previous arts that directly fine-tune the entire ViT parameters, the proposed fine-tuning strategy effectively preserves abundant knowledge from ImageNet and enables the model to adaptively learn forgery-specific features. Additionally, the designed LoRA layers model the long-range interactions within input faces, while the Convpass Adapter layers effectively highlight local forgery anomalies. To this end, the integration of the designed LoRA and Adapter layers leverages the expressivity of transformers and the local forgery priors of CNNs, leading to enhanced generalizability and robustness. Additionally, we design novel Mixture of Experts (MoE) modules within both LoRA and Adapter layers to scale the model’s capacity using fixed activated parameters. The MoE dynamically selects optimal forgery detection experts for input faces, further enhancing the model’s performance.
@@
  • We innovatively integrate LoRA and Adapter modules with the ViT backbone for face forgery detection. This design enables the designed model to simultaneous mine global and local forgery clues in a parameter-efficient manner.
  • We design novel Mixture-of-Experts (MoE) modules to scale up the model capacity. These modules dynamically select optimal forgery experts for input faces, thereby boosting detection performance. Furthermore, the MoE modules can be seamlessly adapted to other transformer architectures in a plug-and-play fashion.
  • We conduct experiments on six Deepfake datasets and various common perturbations. Our experimental results demonstrate that MoE-FFD achieves state-of-the-art generalizability and robustness. Extensive ablation experiments validate the effectiveness of our proposed MoE learning scheme and the designed components.
In this paper, Section 2 comprehensively reviews previous related literature. Section 3 details the proposed mixture of expert framework for face forgery detection. Section 4 presents extensive experimental results under various settings. Finally, Section 5 concludes the paper and discusses possible future research directions.
Figure 2: Overview of the designed MoE-FFD framework. (a) Overall model structure; (b) Details of MoE-FFD transformer block; (c) Details of the designed MoE Adapter layer; (d) Details of each Adapter expert; (e) Details of the designed MoE LoRA layer; (f) Details of each LoRA expert.
@@ 2 Related Works
In this section, we broadly review existing works on face forgery detection, parameter-efficient fine-tuning, and mixture of experts.
@@ 2.1 Face Forgery Detection
Early attempts at face forgery detection primarily relied on extracting handcrafted features, such as the lack of eye blinking [2], inconsistency of head pose [3], face warping artifacts [28], and heart rate anomalies [29]. However, these methods suffer from limited accuracy due to their narrow focus on specific statistical information. In response, learning-based approaches have emerged, leveraging generic network architectures like Xception Network [9], EfficientNet [10], and Capsule Network [30] for forgery feature extraction. Nonetheless, CNN-based methods are prone to overfitting to the training data, resulting in limited generalizability and robustness. Follow-up works, such as Face X-ray [4], F3Net [18], SBI [31], DCL [32], and RECCE [33], have introduced comprehensive forgery frameworks and robust feature extraction techniques, enhancing the model’s generalization capability. With the explosive development of ViT [19], numerous ViT-based methods [20, 21, 22, 23, 24, 25, 26] have been proposed to tackle the Deepfake problem. These approaches take advantage of non-inductive bias and global context understanding, achieving superior face forgery detection performance. However, most of these methods fine-tune the entire ViT model on Deepfake datasets, which is computationally expensive and may lead to loss of valuable knowledge from the ImageNet dataset [34]. In contrast, our approach finetunes external lightweight LoRA and Adapter parameters during training, keeping the ViT backbone fixed with ImageNet weights, thus enabling the model to learn forgery-specific knowledge and enhance the detection performance.
@@ 2.2 Parameter Efficient Fine-Tuning (PEFT)
Over the past few years, the size of deep learning models has exponentially increased, especially after the advent of transformers. Consequently, numerous Parameter Efficient Fine-Tuning (PEFT) methods have been proposed to reduce the computational and storage overhead. PEFT only updates a small portion of the model parameters while freezing the majority of pretrained weights. Adapter [35, 36] is a typical PEFT method, comprising a down-sample layer and an up-sample layer, generally integrated into transformer layers and blocks. Low-Rank Adaptation (LoRA) [37] layer aims at updating two low-rank matrices, significantly reducing the trainable parameters. Zhong et al. [38] found that convolution operation effectively introduces local prior information for image segmentation. Additionally, Visual Prompt tuning (VPT) [39] augments inputs with extra learnable tokens, which can be regarded as learnable pixels to vision transformers. Neural Prompt Search (NOAH) [40] advances further by incorporating Adapter, LoRA, and VPT into vision transformers and optimizing their design through a neural architecture search algorithm. Scale and Shift Feature Modulation (SSF) [41] introduces to scale and shift parameters to modulate visual features during training, achieving comparable performance compared with full fine-tuning. Convpass [42] integrates convolutional bypasses into large ViT models, leveraging local priors to improve image classification performance. However, the application of PEFT methods to face forgery detection remains largely unexplored. In this work, we incorporate dedicated forgery Adapter and LoRA layers to mine global and local clues and achieve a parameter-efficient face forgery detection.
@@ 2.3 Mixture of Experts (MoE)
Mixture-of-Experts (MoE) [43] models aim to augment the model’s capacity without increasing computational expenses. MoE comprises multiple sub-experts and incorporates a gating mechanism to dynamically select the most relevant Top-k experts, thereby optimizing the results. The concept of MoE has been widely used in both Computer Vision [44, 45, 46, 47] and Natural Language Processing [48, 49, 50]. Sparse MoE [48] introduces a router to select a subset of experts, ensuring that the inference time is on par with the standalone counterpart. Subsequent methods such as [51, 52, 53] seek to design novel gating mechanisms to enhance the performance on specific tasks. Follow-up works [51, 54, 55] propose leveraging multi-task learning to guide the model to select optimal experts for a given input query. Moreover, some studies apply MoE architectures to domain adaptation [56] and domain generalization [57] tasks. Previous works usually adopt Feed-Forward Networks (FFN) as the expert choices [48, 44, 58, 59, 60, 61]. However, Feed-Forward Networks (FFNs) still consume significant memory and computational resources during training and inference. In our work, we draw inspiration from the concept of Mixture-of-Experts (MoE) to automate the selection of different LoRA and Adapter modules for various test data. The use of MoE introduces negligible computational overhead, facilitating efficient Deepfake detection. Additionally, the designed MoE modules intelligently select optimal experts, significantly outperforming previous detection methods.
3 Methodology
Herein, we first briefly review the basic preliminaries on Vision Transformer (ViT), Low Rank Adaptation (LoRA), and Compass Adapter. Then we introduce the overview of our designed face forgery detection framework. Finally, we delve into the details of the designed architectures and objective functions of our method.
@@ 3.1.1 Vision Transformer
Vision Transformer consists of multiple blocks of Self-Attention and Multi-Layer Perceptron (MLP). For a given input sequence $x\in\mathbb{R}^{N_{t}\times D}$, $x$ is firstly projected to queries $Q\in\mathbb{R}^{N_{t}\times dim}$, keys $K\in\mathbb{R}^{N_{t}\times dim}$, and values $V\in\mathbb{R}^{N_{t}\times dim}$ using three learnable matrices $W_{q}\in\mathbb{R}^{D\times dim}$, $W_{k}\in\mathbb{R}^{D\times dim}$, and $W_{v}\in\mathbb{R}^{D\times dim}$, where $N_{t}$, $D$, and $dim$ denote the token number, embedding dimension, and hidden dimension, respectively. The $Q$, $K$, and $V$ are calculated by:
@@
3.1.2 Low Rank Adaptation (LoRA)
LoRA is one popular parameter-efficient tuning method. For a pretrained ImageNet weight matrix $W\in\mathbb{R}^{D\times dim}$, LoRA freezes $W$ during training while adding a product of two trainable low-rank matrices $W^{down}W^{up}$, where $W^{down}\in\mathbb{R}^{D\times r}$, $W^{up}\in\mathbb{R}^{r\times dim}$, and rank $r<<min(D,dim)$. As such, the forward pass is modified as:
$h=x(W+\Delta W)=xW+xW^{down}W^{up},$ (3)
where $W$ could be either Query, Key, or Value matrix, i.e., $W\in{\{W_{q},W_{k},W_{v}\}}$. And $W$ is frozen during training. Compared with previous ViT-based finetuning methods, the proposed LoRA learning scheme significantly reduce the trainable parameters.
@@ 3.1.3 Convpass Adapter
Adapter integrates local priors for visual tasks, constructing Convolutional Bypasses (Convpass) within the Vision Transformer (ViT) framework as adaptation modules. The Convpass Adapter can be formulated as:
@@
where the parameters in MLP are fixed with ImageNet weights during training. Convpass generally consists of several trainable convolutional layers. Compared with the MLP layer, the Convpass layer saves computational resources.
@@ 3.2 Overview of the Proposed Framework
Fig. 2 illustrates the designed MoE-FFD framework. Fig. 2 (a) depicts the ViT backbone, which is initialized with the ImageNet weights. The input face firstly undergoes the patch and position embedding. Then, the input tokens are delivered to the designed transformer blocks. Fig. 2 (b) presents more details regarding the designed transformer blocks. We modify the standard ViT blocks by integrating the designed MoE Adapter layer and MoE LoRA layer with the original blocks.
Figure 3: Illustration of the designed Convpass Adapter experts. (a) Vanilla Convolution, (b) Angular Difference Convolution (ADC), (c) Central Difference Convolution (CDC), (d) Radial Difference Convolution (RDC), and (e) Second-Order Convolution (SOC).
@@
Similarly, the attention weights $W_{q,k,v}$ in the MoE LoRA layer are fixed to the ImageNet weights. As shown in Fig. 2 (e), the MoE LoRA module consists of one gate $G_{L}(\cdot)$ and $N$ experts, each with a unique rank. The gate $G_{L}(\cdot)$ selectively activate sparse experts for detecting face forgeries. Fig. 2 (f) depicts the LoRA structure, where $\otimes$ indicates matrix multiplication. The output feature is derived by multiplying the input with two learnable low-rank matrices. The resulting $x_{out}$ is the element-wise summation of the fixed weights and the learned MoE LoRA weights, which are further processed by a self-attention mechanism.
@@
where $L_{ce}$ represents the cross-entropy loss. Following [48, 62], we apply an additional loss $L_{moe}$ to encourage all experts to have equal importance. Further details on $L_{moe}$ will be dedicated at the end of this section.
@@ 3.3 Details of the MoE
TABLE I: Cross-dataset evaluation on five unseen datasets. ‘*’ indicates the trained model provided by the authors. ‘†’ indicates our re-implementation using the public official code. Methods highlighted in blue denote video-level results. #Params indicates the activated parameter number during training.
Method Venue #Params CDF WDF DFDC-P DFD DFR Avgerage
@@
3.3.2 MoE Adapter Layer.
$x_{out}={\rm MLP}(x_{in})+\sum_{j}^{M}G_{A}(x_{in})_{j}E_{A}(x_{in})_{j},$ (8)
$\rm Conv_{1\times 1}^{down}$ and $\rm Conv_{1\times 1}^{up}$ are two $1\times 1$ convolution layers, which down-sample and up-sample the channels, respectively. $\rm Conv_{3\times 3}^{j}$ indicates the specific convolution layer in different experts. We develop $M=5$ types of convolution. These convolutions aim to model different local interactions, facilitating the model to capture abundant local forgery features from input faces.
Fig. 3 shows the designed five convolutions, including (a) Vanilla Convolution, (b) Angular Difference Convolution (ADC), (c) Central Difference Convolution (CDC), (d) Radial Difference Convolution (RDC), and (e) Second-Order Convolution (SOC). The red arrow $\rightarrow$ indicates the subtraction operation. We first formulate the vanilla convolution as (we omit the bias for concision):
$y=\sum_{p\in\Omega_{1}}w_{p}x_{p},$ (10)
3.3.3 Gating Network.
$G(x)={\rm Softmax}({\rm Topk}(H(x),k)).$ (13)
$L_{moe}=CV(Importance(X))^{2};~{}Importance(X)=\sum_{x\in X}G(x),$ (14)
where $CV(\cdot)$ indicates the Coefficient of Variation:
$CV(Importance(X))=\frac{{\rm Mean}[Importance(X)]}{{\rm Std}[Importance(X)]}.$ (15)
As such, the additional loss $L_{moe}$ encourages all experts to have equal importance.
4 Experiments
In this section, we evaluate our model in terms of generalizability, robustness, and parameter-efficiency under a wide variety of experimental settings. We conduct extensive ablation studies to demonstrate the effectiveness of the designed network architecture and the adopted training strategy.
4.1 Implementation Details
We apply the popular MTCNN face detector [68] to crop the face regions. The proposed framework is implemented on the Pytorch [69] platform. The model is trained using Adam optimizer [70] with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999. We set the learning rate of gating network and other trainable parameters as 1e-4 and 3e-5, respectively. The loss weigh $\lambda$ in Eq. (5) is set as 1. We train the model for 20 epochs on one single 3090 GPU with batch size 32. We follow the official dataset split strategy in [6] for fair comparison.
TABLE II: Cross-manipulation detection AUC on the unseen manipulation technique (FF++ C23 dataset).